How do I answer this question?
The odds against WildHorse winning the third race are 11:2. If Molly places a $4 bet on WildHorse to win and WildHorse wins, find Molly’s net winnings
A) $5.50 B) $44 C) $16.50 D) $11 E) $22

Answer by stanbon(48568)

The odds against WildHorse winning the third race are 11:2. If Molly places a $4 bet on WildHorse to win and WildHorse wins, find Molly’s net winnings
A) $5.50 B) $44 C) $16.50 D) $11 E) $22
----------------------------
$11 is paid for every $2 bet
EQIATOPM"
x/4 = 11/2
x = 44/2 = $22
=======================
Cheers,
Stan H.